OpenStudy (baby456):

algebra 1 help 2x+y=10 solve for x Plesae Help me

OpenStudy (anonymous):

2x +y = 10 2x = 10 -y x = 5 - 1/2y

OpenStudy (anonymous):

make the equation to where y is on one side of the equal sign and x is on the other side